im suppose to write an equation of the form y=mx+b for a graph model. one point on the graph model is National CITY center (number of stories) 23, (height) 301.
Answer by scott8148(6628)

since zero stories corresponds to zero height , the graph should probably pass through the origin (0,0)

h = (301/23)s + 0 ___ or ___ h = (301/23)s